Last month, you lent a work colleague $5000 to cover some overdue bills. He agreed to pay you in 1 month with interest at 2% for the month, thus owing you $5100. Today, when the repayment is due, he asked you to extend the loan for another month and he would pay you the $5100 next month. In the meantime, you have had the offer to invest as much as you wish in an oil-well venture that is expected to pay 40% per year and a hot new IT stock that is estimated to return 45% the first year. If you let your colleague have another month, what is the opportunity cost of your decision? (Note: Express your answer in dollar and percentage amounts.) The opportunity cost is $ The opportunity cost in percentage is 1%.
Answer #1

a) Opportunity cost

= 5100*45%*(1/12)

= (5100*0.45)/12

= $191.25

b) Opportunity cost (in percentage term)

= 191.25/5100

= 0.0375 i.e, 3.75%